### v2

# DefenderBench: A Toolkit for Evaluating Language Agents in Cybersecurity Environments

Abstract
Large language model (LLM) agents have shown impressive capabilities in human language comprehension and reasoning, yet their potential in cybersecurity remains underexplored. We introduce DefenderBench, a practical, open-source toolkit for evaluating language agents across offense, defense, and cybersecurity knowledge-based tasks. DefenderBench includes environments for network intrusion, malicious content detection, code vulnerability analysis, and cybersecurity knowledge assessment. It is intentionally designed to be affordable and easily accessible for researchers while providing fair and rigorous assessment. We benchmark several state-of-the-art (SoTA) and popular LLMs, including both open- and closed-weight models, using a standardized agentic framework. Our results show that Claude-3.7-sonnet performs best with a DefenderBench score of 81.65, followed by Claude-3.7-sonnet-think with 78.40, while the best open-weight model, Llama 3.3 70B, is not far behind with a DefenderBench score of 71.81. DefenderBench’s modular design allows seamless integration of custom LLMs and tasks, promoting reproducibility and fair comparisons. An anonymized version of DefenderBench is available at https://github.com/microsoft/DefenderBench.
1 Introduction
LLMs (Touvron et al., 2023a, b; OpenAI, 2023) have demonstrated impressive capacities for understanding and generating natural language. To better leverage LLMs for real-world problem-solving, recent works (Zhao et al., 2024; Park et al., 2023; Wang et al., 2023; Wu et al., 2024a) have integrated LLMs into agentic frameworks, enabling them to perform tasks by interacting with an environment (ecosystem), communicating with multiple agents, and breaking down complex tasks into simpler ones to achieve a higher degree of automation. Recent studies have shown that LLM-based agentic systems effectively handle diverse tasks such as software development (Qian et al., 2024), document-level machine translation (Wu et al., 2024b), and fact-checking (Du et al., 2024). Several concurrent studies have introduced evaluation benchmarks to better assess the capabilities of LLM-based agentic systems, including AgentBench (Liu et al., 2024a) for system and database operations, MLAgentBench (Huang et al., 2024) for machine learning research, SWE-bench (Jimenez et al., 2024) for software development, SmartPlay (Wu et al., 2024c) for games, and WebArena (Zhou et al., 2024) for web workflows. However, how LLM-based agents address cybersecurity-related tasks remains underexplored. Although some contemporaneous works have begun developing evaluation benchmarks for LLM agents in cybersecurity, such as Cybench (Zhang et al., 2024a) for Capture The Flag challenges, CyberMetric (Tihanyi et al., 2024) for cybersecurity knowledge question answering, and CyberSecEval (Bhatt et al., 2024) for code vulnerability detection and exploitation, they focus solely on one or a few specific cybersecurity tasks.
To further explore the capabilities of LLM agents in cybersecurity and enhance fairness of model comparisons and reproducibility, we introduce DefenderBench, a toolkit for evaluating LLM-based agents on cybersecurity tasks. As a dual-use technology (Zhang et al., 2024a; Biden, 2023), LLM agents for cybersecurity are evaluated on three types of tasks: offense, defense, and cybersecurity knowledge understanding. For offense tasks, we implement a text-based wrapper around a network intrusion environment with various configurations. For defense tasks, we include malicious content detection, code vulnerability detection, and code vulnerability fixing. Additionally, we incorporate a multiple-choice question-answering task to assess LLM agents’ understanding of cybersecurity knowledge. Inspired by existing LLM agentic frameworks (Wu et al., 2024c; Liu et al., 2024a; Wei et al., 2022), we introduce an agent baseline to benchmark different LLMs on these cybersecurity tasks. We evaluate several LLMs including open-weight models from the Llama (Dubey et al., 2024) and Phi (Abdin et al., 2024) families, along with proprietary models such as the GPTs (OpenAI, 2023) and Claudes [^1] https://www.anthropic.com/claude. Our experiments show that Claude-3.7-sonnet is the best-performing LLM with a DefenderBench score of 81.65.
To summarize, the contributions of this paper are as follows:
  1. 1.
    We develop an open-source toolkit, DefenderBench, for evaluating LLM-based agents on interactive cybersecurity tasks. This toolkit streamlines data preparation and model evaluation procedures, ensuring fair comparisons. We responsibly release DefenderBench with our benchmark for research purposes.
  2. 2.
    DefenderBench is highly modular, allowing users to easily integrate their own LLMs and agents, as well as add new tasks through a plugin system.
  3. 3.
    We establish a baseline agent and evaluate a wide range of LLMs using DefenderBench, providing a comprehensive assessment of their capabilities in cybersecurity tasks.
2 Related Work
LLM for Cybersecurity. With our growing reliance on digital and interconnected systems and the increasing sophistication of cyber threats (Thakur et al., 2015), cybersecurity has become a critical area of focus. Cybersecurity encompasses a comprehensive range of practices, tools, and strategies aimed at protecting computer systems, networks, and data from unauthorized access, attacks, damage, or disruptions (Li and Liu, 2021; Zhang et al., 2024c). Traditional cybersecurity approaches, such as rule-based systems, struggle to keep pace with rapidly evolving cyber threats. With advancements in LLMs, efforts have been made to leverage LLMs to address cybersecurity challenges. For instance, domain-specific datasets have been curated to fine-tune LLMs for tasks such as program repair (Silva et al., 2023), cybersecurity training (Zhang et al., 2023), network security (Rigaki et al., 2024) and secure code generation (Mechri et al., 2025). Additionally, LLM agents have been employed in tasks like website hacking (Fang et al., 2024b), code vulnerability exploitation (Fang et al., 2024a), debugging (Lee et al., 2024), and penetration testing (Deng et al., 2023). In this paper, we focus on developing a standardized toolkits for evaluating LLM agents.
LLM Agent Benchmark. To evaluate the capabilities of LLM agents, several benchmarks have been developed. AgentBench (Liu et al., 2024a) assesses LLMs across five diverse environments, including operating systems and databases, to evaluate reasoning and decision-making abilities. MLAgentBench (Huang et al., 2024) focuses on machine learning experimentation tasks, testing agents on tasks ranging from improving model performance to addressing research problems. SWE-bench (Jimenez et al., 2024) evaluates LLMs on real-world software issues sourced from GitHub, requiring models to generate patches that resolve described problems. SmartPlay (Wu et al., 2024c) introduces a suite of games to test various capabilities of LLMs, such as planning and spatial reasoning. WebArena (Zhou et al., 2024) provides a realistic web environment for building autonomous agents, enabling the assessment of LLMs in web-based tasks.
Cybersecurity-Specific Benchmarks. In the cybersecurity domain, specialized benchmarks have been introduced. Cybench (Zhang et al., 2024a) offers a framework for evaluating LLM agents on 40 professional-level Capture The Flag (CTF) tasks, encompassing a range of difficulties and scenarios. CyberMetric (Tihanyi et al., 2024) presents a benchmark dataset based on retrieval-augmented generation to assess LLMs’ cybersecurity knowledge. SecEval (Li et al., 2023) provides over 2,000 multiple-choice questions across various cybersecurity domains to evaluate foundation models’ knowledge. CyberSecEval (Bhatt et al., 2024) focuses on code vulnerability detection and exploitation, offering a comprehensive suite for assessing LLMs in secure coding tasks. These benchmarks facilitate targeted evaluations of LLMs in cybersecurity contexts. The closest work to ours is CyberBench Liu et al. (2024b), a benchmark focusing on Natural Language Processing (NLP) tasks related to cybersecurity.
DefenderBench. We introduce DefenderBench, a toolkit designed to evaluate LLM agents in interactive cybersecurity environments. Unlike existing benchmarks mentioned above that focus on specific tasks or domains, DefenderBench encompasses a broad range of cybersecurity-related tasks, covering offense, defense, and knowledge understanding. By integrating insights from general agent benchmarks and adversarial evaluation frameworks, DefenderBench aims to provide a comprehensive assessment platform for LLMs in cybersecurity contexts.
3 Dataset
We describe the datasets included in our benchmark and the preprocessing steps. Currently, DefenderBench consists of five cybersecurity task types.
3.1 Computer Network Intrusion Simulation
In order to protect computer networks against attacks, many organizations conduct red-team network intrustion to proactively detect and remediate vulnerabilites before attackers do. We leverage the network intrusion simulation tool CyberBattleSim (CBS) (Team., 2021) to evaluate the ability of LLM agents to identify vulnerabilities in a network. CyberBattleSim is parameterized by a fixed topology and a set of node vulnerabilities that agents can exploit to move laterally within the network. The goal of the attacker is to take ownership of the network by exploiting vulnerabilities in the computer nodes. We convert CyberBattleSim into a text-based game (Côté et al., 2019) which describes the currently discovered network as some structured text (i.e., JSON) and provides textual feedback in response to the agent’s actions. There are three action types for an attacker to interact with the network:
  • local_vulnerability [src] [type]  # Local exploit (e.g., search credentials in bash history).
  • remote_vulnerability [src] [target] [type]  # Remote exploit (e.g. browse parent directory).
  • connect [src] [target] [port] [credential]  # Connects to a node using leaked credentials.
where [src] refers to the node from which to execute the action, [target] is the node to be exploited, [type] is the type of attack, and [port] is the port used to connect to the target node with the right [credential]. We follow the original CyberBattleSim’s implementation and evaluate on two type of network configurations: a chain network (CBS-Chain) and a capture the flag (CBS-CTF). We report the winning rate (i.e., the number of nodes taken over by the agent divided by the total number of nodes in the network) as the metric for this task.
3.2 Malicious Content Detection
Malicious-Text: for this task, we utilize the dataset processed by Alvarado (2024). [^2] https://huggingface.co/datasets/ealvaradob/phishing-dataset This dataset incorporates two data sources, namely email and text messages, for malicious content detection. The entire dataset contains 20,137 samples labeled as $\{malicious,legitimate\}$. We follow the split of Alvarado (2024), using 80% of the data as the training set and 20% as the test set. To reduce the cost of performing LLMs on our benchmark, we further randomly select 500 samples from the test split as our official test set in the benchmark. Additionally, we select 10 samples per class as the few-shot sampling pool for in-context learning (ICL) Brown et al. (2020). The metric used is the macro-F1 score.
Malicious-Web: This task assesses the ability of LLM agents to discriminate phishing from benign web sites. We use the Phishing Websites Dataset (Ariyadasa et al., 2021) as preprocessed by Alvarado (2024) for malicious website detection. We also discard 144 samples which contain less than 100 characters as they are mostly outliers (e.g. page failed to load). The resulting dataset (15,612 samples) includes 10,220 labeled as legitimate and 5,392 as malicious. We follow the same 80%-20% split as Alvarado (2024) and further uniformly subsample 500 test samples as our test set and 10 training samples per class as the few-shot sampling pool. We report the macro-F1 score for this task.
3.3 Cyber Threat Intelligence (CTI)
MCQA: This task assesses the ability of an LLM agent to understand recent threat intelligence and apply it to challenging questions. A multiple-choice question answering task that uses the CTI-MCQA dataset introduced by Alam et al. (2024). This dataset originally contains 2,500 questions, each associated with a CTI-related webpage or document. After filtering out questions linked to inaccessible webpage or document, we obtained 2,338 samples. We then randomly downsample and split these into a test set (500 questions) and a few-shot sampling pool (20 samples). Each question has four options, with only one correct answer. The metric for this task is macro-F1.
3.4 Code Vulnerability Detection
Vulnerable-CG: This task assesses the ability of LLM agents to detect vulnerabilities in code. We use the code vulnerability detection dataset included in CodeXGLUE (Lu et al., 2021), which is split into training (21,854 samples), validation (2,732 samples), and test sets (2,732 samples). Each sample is a C language function annotated with the label ‘vulnerable’ or ‘non-vulnerable’. Our test samples are 500 randomly selected samples from their test set. We also provide 10 training samples per class as the few-shot sampling pool. The agent’s performance is reported using the macro-F1 score.
Vulnerable-DV: we also include the Devign (Zhou et al., 2019) dataset for code vulnerability detection in our benchmark. Zhou et al. (2019) released two projects, FFmpeg and Qemu, comprising a total of 27,318 samples. We randomly sample 500 samples for our test set. Similarly, we include 10 training samples per class as the few-shot sampling pool and report the macro-F1 score as the evaluation metric.
3.5 Code Vulnerability Fixing
CVEFix: we use the CVEFix dataset (Bhandari et al., 2021) for the vulnerability fixing task. The original dataset contains 12,107 vulnerability fixing commits across 4,249 open-source projects. The dataset includes the source code before and after the changes. We only extract commits with the following conditions: (a) single method modification; (b) the commit is associated to a single CVE (Common Vulnerabilities and Exposures); (c) the programming languages is either: C, C++, Go, Java, JavaScript, PHP, Python, or Rust. As a result, we obtained 240 samples. We use all the samples as the test set for our benchmark. For this task, we provide the method’s source code before the commit and ask the agent to generate a new method that fixes any vulnerability. We report the CodeBLEU score (Ren et al., 2020) between the generated method and the method after the commit.
4 DefenderBench Implementation
4.1 Modules
As depicted in Figure 1, DefenderBench leverages publicly accessible cybersecurity datasets and turns them into interactive environments to evaluate LLM agents. The toolkit comprises three main modules: data preprocessing, task environment, and agent interface. Additionally, we provide instructions to enable users to modify and expand each module.
Figure 1: DefenderBench includes three main modules: data preprocessing, interactive task environment, and agent interface. Along with the framework, we release a baseline agent to evaluate LLMs.
Table 1: Instructions for tasks in DefenderBench.
Task Task Instruction Prompt
CyberBattleSim We are running a network intrusion simulation. Your goal is to take over all the nodes in a network as few steps as possible. As a network attacker, you need to choose one action from these three actions: - local_vulnerability [node_name] [local_attack_type] - remote_vulnerability [source_node] [target_node] [remote_attack_type] - connect [source_node] [target_node] [port] [credential] where [node_name] is the name of the node you want to exploit, ……
Malicious-text You need to carefully analyze whether a given message is malicious or legitimate. The malicious content may be phishing email/SMS, smishing or spam email/SMS. Your answer must be a JSON dictionary like {"answer": "malicious"/"legitimate"} where the value must be either "malicious" or "legitimate”.
Malicious-web You need to carefully analyze whether a given webpage is malicious or legitimate. Your answer must be a JSON dictionary like {"answer": "malicious"/"legitimate"} where the value must be either "malicious" or "legitimate".
MCQA You are given a multiple-choice question (MCQ) from a Cyber Threat Intelligence (CTI) knowledge benchmark dataset. Your task is to choose the best option among the four provided. Your answer must be a JSON dictionary like {"answer": "A/B/C/D"} where the value must be a single letter: A, B, C, or D.
Vulnerability Detection You need to carefully analyze whether a given source code has vulnerability or not. Your answer must be a JSON dictionary like {"answer": "vulnerable"/"non-vulnerable"} where the value must be "vulnerable" or "non-vulnerable".
Vulnerability Fixing You need to carefully analyze a given snippet code and fix its vulnerability. Your answer must be a markdown code block of the same snippet of code once fixed including any existing comments.
Data Preprocessing. The DefenderBench toolkit automatically downloads the required datasets from their respective sources, shuffles the samples randomly according to a fixed random seed, and splits them into a test set and a few-shot sample pool for in-context learning. Once preprocessed, the datasets are cached locally. For network intrusion simulation, we install CyberBattleSim (Team., 2021) as a dependency.
Task Environment. For each task, we set up a task environment that provides task-specific instructions (shown in Table 1), defines the action space for the agent, loads the relevant datasets and constructs few-shot examples if few-shot in-context learning is being conducted (more on this in section 5.3). For the detection, MCQA, and code-fixing tasks, each episode involves presenting the agent with a test sample. Each episode can run for up to five steps. If the agent fails to respond with the expected format, a feedback message is provided and the agent can try again until the episode ends. For the network intrusion task, each episode begins with an initialized network and can run for up to 100 steps to compromise the entire network. The LLM agent interacts with the task environment by providing a text action and the environment provides an observation in return. The observation describes the result of the given action and indicates whether the task has been completed. Additionally, the environment maintains a history of the actions taken by the agent and the corresponding feedback. The history can be provided to the agent as part of its context.
Agent Interface. Our DefenderBench is equipped with an LLM agent interface that enables users to integrate both open- and closed-weight LLMs. Users can also seamlessly incorporate their own agentic system to perform the tasks.
Execution. To evaluate LLM agents on DefenderBench, users can install our toolkit as a Python library. Through a terminal command, users can run all tasks or specify a particular task by using its shorthand name. Additionally, users can choose which LLM to use for the baseline agent. We have also integrated the Weights and Biases library into DefenderBench, [^3] https://wandb.ai/ enabling users to track and visualize their results seamlessly.
Metrics. We report on each task using its original metric as described in Section 3. Inspired by previous evaluation benchmarks like GLUE (Wang et al., 2019), we define a global metric called DefenderBench score, which represents the unweighted average of all task-specific metrics. The DefenderBench score provides an overall indication of performance on cybersecurity tasks.
Baseline Agent. To evaluate the out-of-the-box capability of LLMs in solving cybersecurity tasks, we experiment with a baseline agent with minimal scaffolding in this paper. As illustrated in Figure 1, we begin by providing to the agent a task instruction that explains the task, specifies the response format, and defines the action space. Table 1 shows the task instructions. At each step, the agent is given the trajectory of its prior actions along with the corresponding observations from the environment. At each step, the agent is asked to produce an action in the required format, which is then sent to the task environment to obtain an action observation. Based on this observation, we determine whether the episode should be terminated. If the episode continues, the observation is added to the system prompt as part of the historical trajectory.
5 Experiments
5.1 Backbone LLMs
In our experiments, we use a variety of LLMs as the backbone of our agent. These include (1) open-weight models (Llama 3.1 (Dubey et al., 2024), Llama 3.2, Llama 3.3, and Phi-3.5 (Abdin et al., 2024)), (2) proprietary models (GPT-3.5, GPT-4-turbo, GPT-4o, GPT-4o-mini, Claude-3.5-haiku, and Claude-3.5-sonnet, Claude-3.7-sonnet), and (3) proprietary reasoning models (o1, o1-mini, o3, o4-mini, GPT-4.1, GPT-4.1-mini, and GPT-4.1-nano, Claude-3.7-sonnet-think).
Table 2: DefenderBench test results. CBS: CyberBattleSim, Mal.: Malicious, Vuln.: Vulnerability tasks, CodeBL: CodeBLEU, DefB: unweighted average DefenderBench score.
CBS-Chain CBS-CTF Mal. Text Mal. Web MCQA Vuln.-CG Vuln.-DV CVEfix DefB
win % win % Mac-F1 Mac-F1 Mac-F1 Mac-F1 Mac-F1 CodeBL
Naive Baseline 19.44 22.22 52.40 50.40 25.00 50.00 47.80 83.24 43.81
Open-weight
Llama 3.1 8B 23.61 16.67 88.00 77.20 60.60 49.60 48.60 73.63 54.74
Llama 3.1 70B 77.78 44.44 96.80 83.00 69.80 50.60 51.40 75.88 68.71
Llama 3.2 1B 8.33 16.67 42.00 30.00 50.60 48.60 43.80 66.69 38.34
Llama 3.2 3B 9.72 16.67 83.40 67.00 58.40 46.60 46.40 73.23 50.18
Llama 3.3 70B 100.00 33.33 96.00 82.80 69.60 58.00 57.40 77.31 71.81
Phi-3.5-mini (4B) 8.33 16.67 87.00 66.80 71.00 45.00 44.20 71.97 51.37
Proprietary
GPT-3.5 16.67 16.67 94.20 85.80 61.20 48.00 47.00 54.34 52.99
GPT-4-turbo 90.00 46.67 93.40 83.20 73.80 58.20 57.60 73.72 72.07
GPT-4o 62.50 50.00 93.60 90.00 72.00 55.00 55.20 77.88 69.52
GPT-4o-mini 22.22 19.44 91.40 88.80 67.80 47.60 47.00 79.71 58.00
GPT-4.1 66.67 66.70 89.40 89.80 73.60 19.40 50.60 54.80 63.90
GPT-4.1-mini 50.00 50.00 90.60 89.20 73.60 19.80 45.00 52.80 58.90
GPT-4.1-nano 16.67 16.67 87.00 73.80 63.60 30.00 43.80 48.80 47.50
Claude-3.5-haiku 45.00 40.00 82.70 84.80 67.60 55.20 56.40 70.64 62.79
Claude-3.5-sonnet 100.0 56.67 93.80 88.20 72.40 56.40 56.80 75.74 75.00
Claude-3.7-sonnet 100.0 100.0 96.20 90.00 74.20 56.60 56.00 80.18 81.65
Proprietary reasoning
o1-preview 16.67 16.60 82.50 88.70 77.40 56.40 51.40 50.10 59.70
o1-mini 50.00 50.00 80.30 74.40 37.40 49.60 48.60 53.70 60.30
o3 83.30 20.00 92.40 88.00 76.40 30.80 59.60 55.60 63.90
o4-mini 66.70 20.00 92.00 84.60 70.00 32.20 57.40 52.40 50.80
Claude-3.7-sonnet-tk 100.0 76.67 94.40 91.00 78.20 54.60 52.80 79.50 78.40
5.2 Main Results
For comparison, we included a naive baseline agent. This baseline randomly selects actions from the action list for all tasks except CVEFix. For CVEFix, the naive baseline is a copy-paste agent that outputs the original code without any modifications. We run each evaluation experiment five times and report the average performance in Table 2.
Overall Performance. Claude-3.7-sonnet achieves the highest DefenderBench score of 81.65 across all tasks. Among the open-weight models, the Llama 3.3 70B model attains the highest score of 71.81, outperforming GPT-3.5, which records a score of 52.99. Among the reasoning-focused models evaluated, Claude-3.7-sonnet-think achieves the best performance with a DefenderBench score of 78.40. Comparing overall results, we observe that reasoning-augmented models do not outperform their counterparts on cybersecurity tasks. When comparing models of different sizes, we observe that larger models generally perform better. For example, the 70B version of Llama 3.1 surpasses its 8B variant by 13.97 points, and the 3B-sized Llama 3.2 outperforms its 1B counterpart by 11.84 points. Similarly, GPT-4.1, GPT-4.1-mini, and GPT-4.1-nano achieve scores of 63.90, 58.90, and 47.50, respectively, reflecting a steady decline as model size decreases. As expected, these results highlight the substantial impact of model size on task performance.
Network Intrusion. Network Intrusion. For the CyberBattleSim network intrusion task, LLaMA 3.3 70B, Claude-3.5-sonnet, Claude-3.7-sonnet, and Claude-3.7-sonnet-think achieve a perfect 100% winning rate on the chain-pattern network, successfully compromising all 12 nodes in all five runs. This demonstrates that advanced LLMs are capable of completing network intrusions when the infection pattern across nodes is regular and predictable. In terms of efficiency, the average number of steps to completion is 26.5 for LLaMA 3.3 70B, 57.3 for Claude-3.5-sonnet, 50.2 for Claude-3.7-sonnet, and 43.4 for Claude-3.7-sonnet-think. Notably, LLaMA 3.3 70B completes the intrusion in as few as 24 steps in three of five trials. In contrast, GPT-3.5 performs significantly worse, with an average winning rate of only 16.67%, managing to infect up to three new nodes across five runs. Smaller models, such as LLaMA 3.2 1B and Phi-3.5-mini, also struggle, each achieving a winning rate of just 8.33% and generally failing to compromise any additional nodes. Performance drops substantially in the more complex CyberBattleSim ToyCTF environment, which features a less regular structure and requires more advanced strategic planning. Claude-3.7-sonnet again achieves the best result, maintaining a 100% winning rate and successfully compromising all nodes in the network. However, it requires an average of 75 steps to complete the intrusion, reflecting the greater difficulty of this environment. Other models perform considerably worse in this setting: GPT-4-turbo and LLaMA 3.1 70B achieve winning rates of only 46.67% and 44.44%, respectively. These results suggest that while most top-tier LLMs can effectively handle structured attack scenarios, their capabilities are still limited in more dynamic or irregular environments.
Malicious Content Detection. On malicious content detection tasks, Llama 3.1 70B achieves the best performance on Malicious-Text, with a Macro-F1 score of 96.80, while Claude-3.7-sonnet-think attains the highest score on Malicious-Web, with a Macro-F1 of 91.00. For Malicious-Text, most proprietary LLMs achieve Macro-F1 scores above 90, indicating strong performance, and most open-weight models also perform well, with scores exceeding 80. However, Llama 3.1 1B performs significantly below expectations, failing to surpass the random baseline on both detection tasks. Its especially poor performance on Malicious-Web is likely due to the long sequence length of the HTML input, which poses a challenge for smaller models with limited context windows and capacity.
Vulnerability Detection. Across both Vulnerable-CG and Vulnerable-DV, most models perform only slightly better than the random baseline, indicating the difficulty of identifying subtle flaws in code with limited context information. GPT-4-turbo achieves the highest scores on Vulnerable-CG, with a Macro-F1 of 58.20, and GPT-o3 performs best on Vulnerable-DV with Macro-F1 of 59.60. Among open-weight models, Llama 3.3 70B performs best, achieving Macro-F1 scores of 58.00 and 57.40 on the respective tasks—closely trailing GPT-4-turbo. These results suggest that, despite their strong general capabilities, current LLMs still struggle to robustly detect security vulnerabilities in code, likely due to the need for precise program understanding and fine-grained reasoning. Improving performance on such tasks may require further domain-specific training or integration with program analysis tools.
Figure 2: Test results of augmented experiments. LM: Llama, CLD-3.7-S: Claude-3.7-sonnet, and CLD-3.7-S-tk: Claude-3.7-sonnet-think.
MCQA. The best-performing LLM on the multiple-choice question-answering task is Claude-3.7-sonnet-think, achieving a Macro-F1 score of 78.20. Among open-weight models, surprisingly, Phi-3.5-mini delivers the strongest results, with a Macro-F1 score of 71.00—despite having only 4B parameters.
Code Fixing. For the CVE code fixing task, none of the LLM agents outperform the copy-paste baseline in terms of CodeBLEU scores. This is primarily due to the minimal modifications required to fix code vulnerabilities in the original script, while CodeBLEU compares the entire generated script with the gold script. Among the models, GPT-4o-mini achieves the highest CodeBLEU score of 79.71. The best-performing open-weight model is Llama 3.1 70B, with a CodeBLEU score of 75.88. In contrast, GPT-3.5 performs poorly, achieving only a CodeBLEU score of 54.34. These results suggest that CodeBLEU may not fully reflect patch quality in cases involving small edits. Our future work should explore alternative evaluation metrics better suited to small, targeted code changes. Nonetheless, larger models still demonstrate relatively better capability in capturing precise code edits.
5.3 Auxiliary Analyses
In this section, we provide additional analyses to investigate how LLM agents perform on cybersecurity tasks when equipped with (1) augmented information and (2) chain-of-thought (CoT) prompting. To be cost friendly, we select representative models to evaluate on a subset of our test set, limiting the number of test samples to 100.
Experiments with Augmented Information. We evaluate the performance of LLMs when augmented information is provided. Figure 2 illustrates the results for the malicious content detection, vulnerability detection, and MCQA tasks. For the Malicious-Text and Vulnerable-DV tasks, we include four samples (two per class) in the system instruction. Due to the long input sequence in the Malicious-Web task, we limit the few-shot in-context learning setup to two samples (one per class). For the CTI-MCQA task, we leverage the CTI-related webpages that were originally used to generate the questions, providing them as context information for the agent to utilize.
Across the four detection tasks, we observe that few-shot in-context learning improves the performance of most LLMs. However, it does not yield better results for Llama 3.2 1B and 3B or Phi-3.5 mini, likely due to their limited capacity to process long sequences. Similarly, incorporating related CTI webpages into the MCQA task significantly boosts the performance of LLM agents. For instance, the agents utilizing the Llama 3.2 3B, GPT-4o mini, and Claude-3.7-sonnet models achieve Macro-F1 improvements of 27.00 and 26.60, and 22.2, respectively. In contrast, the performance of the agent with the Llama 3.2 1B model deteriorates substantially, further highlighting its limited ability to handle long sequences effectively. These findings suggest that augmenting LLM inputs with relevant examples or context can substantially boost performance—especially for larger models with higher capacity. For small models, such augmentation may introduce complexity that overwhelms their limited context windows or generation power, leading to performance drops.
Experiments with CoT Agent. Chain-of-Thought (CoT) prompting  (Wei et al., 2022) is a promising technique that leverages LLM’s reasoning capacity to enhance accuracy in target tasks (Hsieh et al., 2023; Zhang et al., 2024b; Li et al., 2025). Hence, We compare our basic agent with an LLM agent utilizing CoT prompting. For the CoT agent, we include a CoT step before asking the agent to decide on an action. The CoT question is framed as: "What is the best action to take? Let’s think step by step." In Table 3, we group tasks into two categories: (1) interactive tasks, which include two network intrusion environments, and (2) static tasks, comprising the other five environments. Our results show that the CoT agent improves the performance of most LLMs. For the interactive environments, GPT-4o and GPT-3.5 achieve notable improvements in average winning rates, with increases of 17.0 and 9.1, respectively. While the CoT agent does not consistently enhance performance for some LLMs on static tasks, we observe improvements for GPT-3.5 and Llama 3.2 1B, with average score increases of 1.4 and 1.2, respectively. These findings suggest that CoT prompting is particularly effective for interactive, multi-step reasoning tasks, where step-by-step deliberation enables more strategic decision-making.
Table 3: Effect of chain-of-thought prompt agent. The green color indicates that the agent with CoT performs better than the basic agent.
Interactive Static DefenderBench
Base CoT Base CoT Base CoT
Llama 3.1 8B 20.1 22.2 66.3 65.8 54.7 54.9
Llama 3.1 70B 61.1 44.5 71.3 70.6 68.7 64.0
Llama 3.2 1B 12.5 12.5 47.0 48.2 38.3 39.3
Llama 3.2 3B 13.2 15.3 62.5 62.9 50.2 51.0
Phi-3.5 mini 12.5 14.6 64.3 63.1 51.4 50.9
GPT-3.5 16.7 25.8 65.1 66.5 53.0 56.3
GPT-4-turbo 68.3 70.8 73.3 72.8 72.1 72.3
GPT-4o 56.3 73.3 73.9 71.5 69.5 71.9
GPT-4o-mini 20.8 23.6 70.4 71.5 58.0 59.5
6 Conclusion
We introduced DefenderBench, a rigorous evaluation benchmark designed to assess LLM agents on cybersecurity tasks. DefenderBench encompasses five diverse tasks spanning offense, defense, and understanding domains. Its modular design allows for seamless integration of custom LLMs and tasks, promoting reproducibility and fair comparisons.
We benchmarked several state-of-the-art and popular LLMs highlighting the superior performance of models like Claude-3.7-sonnet in various cybersecurity tasks. That said, detecting and fixing code vulnerabilities remain a challenging task for even top tier LLMs. We also observed that few-shot in-context learning improves most LLMs’ performance in detection tasks, but smaller models like Llama 3.2 1B struggle with long sequences, while incorporating CTI webpages boosts performance for some models. Furthermore, the simple CoT agent scaffolding enhances most LLMs’ performance, especially in interactive tasks, with notable improvements for GPT-4o and GPT-3.5.
7 Limitations
Benchmark Construction. DefenderBench currently includes only five cybersecurity-related tasks, which we acknowledge is not exhaustive in covering the breadth of challenges in the domain. Additionally, we do not host the data but instead rely on publicly accessible datasets and environments. We aim to expand this benchmark over time and encourage contributions of new datasets and evaluation metrics from the research community.
Model Selection While we have evaluated DefenderBench on a variety of SOTA models, due to the rapid release of new models by varying providers, the results we share here do not cover additional leading models, such as Gemini (Anil et al., 2023), Mistral (Jiang et al., 2024), or DeepSeek (Guo et al., 2025). We hope that DefenderBench will serve as a foundation for future studies to evaluate a more diverse set of LLMs, enabling a comprehensive understanding of their capabilities in cybersecurity tasks.